The market demand for a type of carpet has been estimated as: P = 40 -


0.25Q, where P is price (GHØ/yard) and Q is rate of sales (hundreds of yards per month). The market supply is expressed as: P = 5.0 + 0.05Q. A typical firm in this market has a total cost function given as: C = 100 - 20.0q + 2.0q2.


Determine the equilibrium market output rate and price.


Determine the output rate for a typical firm.


Determine the rate of profit (or loss) earned by the typical firm


 



1
Expert's answer
2021-04-19T18:51:42-0400

When Demand meets Supply for equilibrium Price and Quantity

"Q_d=Q_s \\

160-4P=20P-100 \\

260=24P"

"P_e=10.83" Equilibrium Price

"Q_e=160-4 \times 10.83 =116.66" Equilibrium Quantity

For Typical Firm P=MC

"10.83=4q-20 \\

4q=30.83 \\

q=7.70"

Profit for typical firm

Profit=Revenue-Cost

"=P \times q-C \\

=10.83 \times 7.70 -100+20 \times 7.70 -2(7.70)^2=1883"